Could you please help me with these 2 question.
1. When MUST we have a proper or an improper fraction? for addition, subtraction, multiplication, or division?
2. Translate in to an algebra equation: 5 less than two thirds of a number is three.

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

2) 2/3x-5=3
2/3x=8
x=8*3/2
x=24/2
x=12
-----------
2/3(12)-5=3

Answer by stanbon(75887)

1. When MUST we have a proper or an improper fraction? for addition, subtraction, multiplication, or division?
addition can be done with either
subtraction can be done with either
Multiplication can be done with either
division can be done with either
--------------------------------------------------------------
2. Translate into an algebraic equation: 5 less than two thirds of a number is three.
Equation:
(2/3)x-5 = 3
========================
Cheers,
Stan H.